OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and has a slope of 5? y = 5x - 3 y = 5x - 1 y = 5x + 3 y = 5x - 19

jimthompson5910 (jim_thompson5910):

which one do you think it is

OpenStudy (anonymous):

3rd one

jimthompson5910 (jim_thompson5910):

y = 5x + 3 1 = 5(4) + 3 ... Plug in x = 4 and y = 1 1 = 20 + 3 1 = 23 which is false

jimthompson5910 (jim_thompson5910):

so it can't be the 3rd one

OpenStudy (anonymous):

ok then help plz

jimthompson5910 (jim_thompson5910):

y = mx+b y = 5x + b ... Plug in the given slope m = 5 1 = 5*(4) + b ... Plug in the given point (4,1) 1 = 20 + b 1 - 20 = b -19 = b b = -19 So the equation is y = 5x - 19

jimthompson5910 (jim_thompson5910):

We can check it y = 5x - 19 1 = 5(4) - 19 1 = 20 - 19 1 = 1 which is true

OpenStudy (anonymous):

yay

jimthompson5910 (jim_thompson5910):

glad it's making sense now

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np